Hydrogen initially occupies 164 liters. If it changes to a volume of 73 L at 0° C and 760 mm Hg. If the initial temperature of the hydrogen was 28° C, what was its initial pressure of the gas in mm Hg? Round to two decimal places.


1
Expert's answer
2021-03-05T03:27:35-0500

0oC = 273.15 K

28oC = 301.15 K

According to the combined gas law,

"\frac{p_1V_1}{T_1}=\frac{p_2V_2}{T_2}"

Therefore,

"p_1=\frac{T_1p_2V_2}{V_1T_2}=\frac{301.15K\times760mmHg\times73L}{164L\times273.15K}=372.97mmHg" (to 2 decimal places)


Answer: 372.97 mmHg